A grocer combined walnuts that cost P700 per kilo and cashew nuts that cost P450 per kilo. Find the amount of walnuts and the amount of cashew nuts used to make the 50 kilo mixture costing P525 per kilo.
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39631)

total cost?
total kilograms?

Ratio is the price.

http://www.algebra.com/my/mixture-price-two-part-both-parts-unknown.lesson?content_action=show_dev

You can do this in one single variable as an alternative. Price is like a concentration, and no matter if price or percents, same idea:
Here is an example: Using just one variable to solve two-part mixture

Answer by ikleyn(52933)

.
A grocer combined walnuts that cost P700 per kilo and cashew nuts that cost P450 per kilo. Find the amount of walnuts and the amount of cashew nuts used to make the 50 kilo mixture costing P525 per kilo.
----------------------------------------------------------------
x kilo of walnuts and y kilo of cashew nuts.

    x +     y = 50,       (1)
700*x + 450*y = 50*525.   (2)

Solve it. For example, use the substitution method.
Express x via y from (1): x = 50-y, and substitute it into (2).
You will get single equation for y.

700*(50-y) + 450*y = 26250.

Solve it. The rest is arithmetic.

Good luck!
Happy new year!